# Probabilistically Tightened Linear Relaxation-based Perturbation Analysis for Neural Network Verification

### v1

@@ 2.2 Linear Relaxation-based Perturbation Analysis (LiRPA) approaches
For a DNN composed of $N$ layers, the base case is defined as $\underline{{\bm{A}}}^{(N)}=I$. Moreover, in the case of a single output node, $\underline{{\bm{A}}}^{(N-1)}={\bm{w}}^{\top}$, where ${\bm{w}}$ is the weight vector of the final layer. For the remaining layers $i\in\{N-2,\dots,1\}$, the linear relaxation is propagated using the recurrence $\underline{{\bm{A}}}^{(i)}=\underline{{\bm{A}}}^{(i+1)}\,\underline{{\bm{D}}}^{(i)}\,{\bm{W}}^{(i)}$, where ${\bm{W}}^{(i)}$ is the inter level weight matrix and $\underline{{\bm{D}}}^{(i)}$ is a diagonal matrix encoding the linear relaxation of the activation function at the $i$-th layer. For each layer $i$, we also recursively compute a bias vector $\underline{{\bm{b}}}^{(i)}$ based on the reachable sets of the $i$-th layer nodes and the matrix $\underline{{\bm{A}}}^{(i+1)}$. Each diagonal entry $\underline{{\bm{D}}}^{(i)}_{j,j}$ depends on the preactivation bounds of neuron $j$ (computed during the forward pass) and the sign of $\underline{{\bm{A}}}_{j}^{(i+1)}$, i.e., the coefficient associated with neuron $j$ in the linear relaxation of the next layer.222 The explicit formulas used for computing $\underline{{\bm{D}}}^{(i)}$ and ${\bm{b}}^{(i)}$ can be found in the discussion of the example provided below. At the end of the process, we obtain a final linear lower bound function given by
${\bm{a}}^{T}_{\text{LiRPA}}({\bm{x}})+{\bm{c}}_{\text{LiRPA}},$
where ${\bm{a}}^{T}_{\text{LiRPA}}({\bm{x}})=\underline{{\bm{A}}}^{(1)}{\bm{x}}$ and ${\bm{c}}_{\text{LiRPA}}=\underline{{\bm{A}}}^{(N)}\underline{{\bm{b}}}^{(N-1)}+\cdots+\underline{{\bm{A}}}^{(2)}\underline{{\bm{b}}}^{(1)}$.
A lower bound on the minimum of $f$ in $\mathcal{C}$ (the $\ell_{\infty}$ norm ball around ${\bm{x}}_{0}$) is then easily obtained using Hölder’s inequality [54] as
$\min\limits_{{\bm{x}}\in\mathcal{C}}{\bm{a}}^{T}_{\text{LiRPA}}({\bm{x}})+{\bm{c}}_{\text{LiRPA}}=-||\underline{{\bm{A}}}^{(1)}||_{1}\cdot\varepsilon+\underline{{\bm{A}}}^{(1)}{\bm{x}}+{\bm{c}}_{\text{LiRPA}}.$ (1)
@@
In the following, we provide a simple example of linear bound computation for a toy DNN depicted in Figure 2 using a state-of-the-art LiRPA-based approach, namely CROWN [54]. We consider a neural network with two inputs, two hidden layers with ReLU activation, and one single output. Following the notation introduced in Sec. 2 we define:
@@
$x_{1}$$x_{2}$$z_{1}^{(1)}$$z_{2}^{(1)}$$\hat{z}_{1}^{(1)}$$\hat{z}_{2}^{(1)}$$z_{1}^{(2)}$$z_{2}^{(2)}$$\hat{z}_{1}^{(2)}$$\hat{z}_{2}^{(2)}$$f(x)$222233-3- 311114444ReLUReLU4444222222-2- 21111ReLUReLU22-2- 21111 $[-2,2]$ $[-1,3]$ ${\bm{z}}^{(1)}$ $[-5,7]$ $[-10,18]$ $[0,18]$ $\hat{{\bm{z}}}^{(1)}$ $[0,7]$ ${\bm{z}}^{(2)}$ $[-36,28]$ $[0,32]$ $\hat{{\bm{z}}}^{(2)}$ $[0,28]$ $[0,32]$ $\bm{[-33.0,18.9]}$ $[-56,32]$ [42.,24.3][-42.,24.3][ - 42 . , 24.3 ] $z^{(3)}$
Figure 2: Toy DNN used in this example. Intervals reported in green are the exact output reachable set computed via MIP, in black are the ones of IBP, and finally, in purple, the results for CROWN considering the input $[[-2,2],[-1,3]]$.
By propagating these intervals through the DNN, we obtain the interval $[-56,32]$ as the output reachable set. Given the reasonable size of the neural network, before computing the linear lower and upper bounds using LiRPA, we employed mixed integer programming (MIP) [42] solution to compute the true min and max of the function, respectively, which correspond to $[-33.0,18.89]$, highlighted in green in Fig. 2.
To compute the lower and upper bounds using CROWN [54], we employ LiRPA’s backward computation strategy. To this end, it is helpful to represent the neural network as reported in Fig. 3.
@@
We note that $\hat{{\bm{z}}}^{(2)}$ and $\hat{{\bm{z}}}^{(1)}$ contain non-linear activation functions (ReLU), and we have to linearize them to keep the linear relationship between the output and these hidden layers. To this end, we create $2\times\#$ReLU layers (for the lower and upper bound, respectively) diagonal matrices $\underline{{\bm{D}}}^{(2)},\overline{{\bm{D}}}^{(2)},\underline{{\bm{D}}}^{(1)},\overline{{\bm{D}}}^{(1)}$ and bias vectors $\underline{{\bm{b}}}^{(2)},\overline{{\bm{b}}}^{(2)},\underline{{\bm{b}}}^{(1)},\overline{{\bm{b}}}^{(1)}$ reflecting the impact of each ReLU nodes on the final output. We report for simplicity here the original definition provided in [54] (a similar definition is applied to compute the $i$-th layer $\overline{{\bm{D}}}^{(i)}$ and $\overline{{\bm{b}}}^{(i)}$):
$\underline{{\bm{D}}}^{(i)}=\begin{cases}1\quad&{\bm{l}}_{j}\geq 0,\\ 0\quad&{\bm{u}}_{j}\leq 0,\\ \alpha_{j}\quad&{\bm{u}}_{j}>0>{\bm{l}}_{j}\text{ and }{\bm{A}}_{j}^{(i+1)}\geq 0,\\ \frac{{\bm{u}}_{j}}{{\bm{u}}_{j}-{\bm{l}}_{j}}\quad&{\bm{u}}_{j}>0>{\bm{l}}_{j}\text{ and }{\bm{A}}_{j}^{(i+1)}<0\end{cases}$ $\underline{{\bm{b}}}^{(i)}=\begin{cases}0\quad&{\bm{l}}_{j}>0\text{ or }{\bm{u}}_{j}\leq 0,\\ 0\quad&{\bm{u}}_{j}>0>{\bm{l}}_{j}\text{ and }{\bm{A}}_{j}^{(i+1)}\geq 0,\\ -\frac{{\bm{u}}_{j}{\bm{l}}_{j}}{{\bm{u}}_{j}-{\bm{l}}_{j}}\quad&{\bm{u}}_{j}>0>{\bm{l}}_{j}\text{ and }{\bm{A}}_{j}^{(i+1)}<0.\end{cases}$
@@
$\geq\underline{{\bm{A}}}^{(3)}\underline{{\bm{D}}}^{(2)}z^{(2)}({\bm{x}})$ computing a linearization for $\text{computing a linearization for }\hat{z}^{(2)}$
$\geq\underbrace{\underline{{\bm{A}}}^{(3)}\underline{{\bm{D}}}^{(2)}{\bm{W}}^{(2)}}_{\underline{{\bm{A}}}^{(2)}}\hat{z}^{(1)}({\bm{x}})$ rewriting $\text{rewriting }{z}^{(2)}={\bm{W}}^{(2)}\hat{z}^{(1)}$
$\geq\underline{{\bm{A}}}^{(2)}\underline{{\bm{D}}}^{(1)}z^{(1)}({\bm{x}})$ computing a linear bound for $\text{computing a linear bound for }\hat{z}^{(1)}$
$\geq\underbrace{\underline{{\bm{A}}}^{(2)}\underline{{\bm{D}}}^{(1)}{\bm{W}}^{(1)}}_{\underline{{\bm{A}}}^{(1)}}({\bm{x}})$ rewriting $\text{rewriting }{z}^{(1)}={\bm{W}}^{(1)}\hat{z}^{(0)}={\bm{W}}^{(1)}({\bm{x}})$
@@
Finally, we compute the sum in the bias vectors $\underline{d}=\underline{{\bm{A}}}^{(3)}\underline{{\bm{b}}}^{(2)}+\underline{{\bm{A}}}^{(2)}\underline{{\bm{b}}}^{(1)}=-35.88$ and $\overline{d}=\overline{{\bm{A}}}^{(3)}\overline{{\bm{b}}}^{(2)}+\overline{{\bm{A}}}^{(2)}\overline{{\bm{b}}}^{(1)}=12.27$.
The final linear relation is thus $\underline{f}({\bm{x}})\geq\underline{{\bm{A}}}^{(1)}({\bm{x}})+\underline{d}$ and $\overline{f}({\bm{x}})\leq\overline{{\bm{A}}}^{(1)}({\bm{x}})+\overline{d}$. Using Hölder’s inequality [54], we obtain
$\underline{f}_{\text{CROWN}}=\min\limits_{{\bm{x}}\in\mathcal{C}}\underline{{\bm{A}}}^{(1)}({\bm{x}})+\underline{d}$ $=-||\underline{{\bm{A}}}^{(1)}||_{1}\cdot\varepsilon+\underline{{\bm{A}}}^{(1)}{\bm{x}}_{0}+\underline{d}$
@@
$\overline{f}_{\text{CROWN}}=\max\limits_{{\bm{x}}\in\mathcal{C}}\overline{{\bm{A}}}^{(1)}({\bm{x}})+\overline{d}$ $=||\underline{{\bm{A}}}^{(1)}||_{1}\cdot\varepsilon+\overline{{\bm{A}}}^{(1)}{\bm{x}}_{0}+\overline{d}$
@@
As we can notice, we obtain a tight over-approximation of the true lower and upper bounds, significantly improving the bounds derived with naive IBP.
2.3 Related Work
Formal Verification.
In recent years, significant research has been dedicated to formal verification and especially to robustness verification [23, 46]. For example, [4] shares with our work the goal of achieving certified robustness for neural networks, but differs significantly in both focus and methodology. Their approach is specifically designed for the time-series domain, leveraging statistical constraints and polynomial transformations to generate adversarial examples and derive robustness guarantees. In contrast, our work addresses general classification tasks and introduces a probabilistically sound verification method based on tight linear relaxation bounds. Other related work that integrates recurrent neural network(RNN)-based policy learning with formal verification is presented in [11]. Specifically, the authors target policy verification under temporal logic constraints in partially observable setups by extracting finite-state controllers from RNNs to enable model checking. In contrast, our work focuses on classification tasks and specifically robustness verification of general feedforward deep neural networks under input perturbations. The most closely related works to our proposal comprise LiRPA-based verification approaches that focus on increasing the quality of linear bounds of the most popular activation functions, such as ReLU, and more general activation functions. More specifically, [51] proposes a framework for deriving and computing near-optimal sound bounds with linear relaxation-based perturbation analysis for neural networks. This framework is the base of all the most famous state-of-the-art formal verification tools such as CROWN [54], $\alpha$-CROWN [52], $\beta$-CROWN [44], GCP-CROWN [55], the top performer on last years VNN-COMP [32, 8]. All these approaches will be used as a worst-case verification result in our experiments.
Different approaches have tried to incorporate a sampling-based approach to enhance either the linear relaxation of arbitrary non-linear functions [34, 6, 53] or the verification process [3], but still maintaining provable verification certificates. For instance, [6, 34] proposed a method synthesizing linear bounds for arbitrary complex activation functions, such as GeLU [20] and Swish [37], by combining a sampling technique with an LP solver to synthesize candidate lower and upper bound coefficients and then certifying the final result via SMT solvers [18]. In [53], the authors focus on estimating the actual domain of an activation function by combining Monte Carlo simulation and gradient descent methods to compute an underestimated domain, which is then paired with over-approximations to define provable linear bounds. If, on the one hand, our PT-LiRPA also employs a similar sampling-based procedure to compute the estimated domain of reachable sets, on the other hand, it differs fundamentally in both nature and focus. Specifically, our approach provides an explicit formula that, for any desired level of confidence, gives the number of samples sufficient to estimate the minimum and maximum pre-activation value of any node (over the set of inputs in $\mathcal{C}$) with the given confidence. In addition, we are also able to estimate the maximum error between our estimates and the true extremal pre-activation values. Crucially, our approach adopts a probabilistic perspective to derive safety insights, whereas [53] relies on combining under- and over-approximations to produce provable bounds—an approach that may inherit the limitations of formal methods discussed earlier.
Probabilistic Verification.
Recently, several works have explored probabilistic verification of machine learning models. For example, [16, 31, 56] focus on a probabilistic verification perspective for general machine learning models (e.g., deep generative/diffusion models), leveraging either uncertainty sources generated by specific encoders or using symbolic reasoning and probabilistic inference. [45] proposes a Monte Carlo-based method utilizing multi-level splitting to estimate the probability of rare events for robustness verification. Other approaches rely on Chernoff-Cramér bounds, e.g., [14] estimates the local variation of neural network mappings at training points to regularize the loss function. In [26], the authors probabilistically certify a neural network by overapproximating input regions where robustness is violated. In [35] a probabilistic certification approach is proposed that can be used in general attack settings to estimate the probability of a model failing if the attack is sampled from a certain distribution. Another line of work, like [49] uses extreme value theory (EVT) [19] to statistically estimate the local Lipschitz constant and assess the probabilistic robustness of the model without relying on the overapproximation of LiRPA-based solvers. Finally, recent works [5, 29, 30, 39, 7] focus either on different types of verification, such as probabilistic enumeration of (un)safe region for neural networks, or on probabilistic specifications for robustness verification, which falls outside the scope of this paper, as we are interested in standard robustness verification with common specifications.
More closely related to our work are the approaches used in PROVEN [48] and the Randomized Smoothing of [12], which also focus on probabilistic robustness guarantees for neural network classifiers. PROVEN builds upon LiRPA-based techniques, combining them with concentration inequalities to derive probabilistic bounds on the network’s output under input perturbations. Similarly, Randomized Smoothing constructs a smoothed classifier by adding random noise, typically Gaussian, to the input and then provides robustness guarantees by analyzing the output distribution of the smoothed model. PT-LiRPA aligns with these approaches in that it also leverages linear relaxation techniques and probabilistic reasoning, but introduces a novel perspective by focusing on how much sampling-based underestimations affect the linear bounds used during network linearization. This enables a finer-grained analysis of robustness with strong probabilistic guarantees, which is particularly useful in scenarios where the available methods for computing exact worst-case bounds either fail (by not terminating under reasonable time and space constraints) or terminate with only over-conservative bounds.
@@ 3 Probabilistically Tightened LiRPA via Underestimation
In this section, we present all the theoretical and practical components to compute tightened bounds with probabilistic guarantees within our PT-LiRPA approach. Our approach is based on a statistical methodology that allows to compute estimates on the neural network’s output in the form $\min_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})$ and $\max_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})$, which hold with high confidence. In particular, we employ the tools of statistical prediction of tolerance limits of [50]. Fix a node $z^{(i)}_{j}$ in the neural network and, as before, let $z^{(i)}_{j}({\bm{x}})$ be the preactivation value for the node when the input to the network is a vector ${\bm{x}}.$ Let ${\bm{x}}_{1},\dots\bm{x}_{n}$ be $n$ points (vectors) independently and uniformly sampled from the perturbation set of interest $\mathcal{C}.$ We compute $z^{(i)}_{j}$’s (estimated) pre-activated bounds as:
@@
However, based on the results of [50], for any $\psi,R\in(0,1)$, we can choose the sample size $n$ that guarantees that the estimated reachable set is correct with probability $\psi$ for at least a fraction $R$ of points in the perturbation set $\mathcal{C}$. Crucially, this statistical result does not require any knowledge of the probability distribution governing the output of our function of interest. Formally, we have the following.
@@
[…]
Hence, following Lemma 3.1, whose proof follows from [50], for any desired confidence level $\psi$, and lower bound fraction $R$, we can compute the number $n$ of samples sufficient to obtain the provable probabilistic guarantees on the desired reachable set accuracy. Specifically, we have that if we use a sample of $n\geq\frac{\ln{(1-\psi)}}{\ln{(R)}}$ input points and with them we obtain an estimated reachable set $[\overline{l}^{(i)}_{j},\underline{u}^{(i)}_{j}]$, then with probability $\psi$ at most a fraction $(1-R)$ of points in an indefinitely larger future sample could fall outside such reachable set. By taking into account the total number of neurons in the DNN, and using a independently chosen set of points for each neuron, we can extend the above result to compute reachable sets for each neuron that are simultaneously correct with probability $\psi$ for at least a fraction $R$ of the points in $\mathcal{C}$:
@@
We will now show that we can combine the (probabilistically valid) estimation computed on the reachable sets of individual neurons using a LiRPA approach, so as to obtain, first, a probabilistically valid over-approximation of any ReLU layer in the DNN and thus on the final network’s lower and upper outputs. In detail, we start proving that the combination of the estimated reachable sets and the linearization approach employed for any (ReLU) layer of the networks results in a probabilistic valid over-approximation for at least a fraction $R$ of the perturbation region $\mathcal{C}$.
@@
Consider a ReLU DNN with $m$ neurons distributed over $N$ layers, and $\mathcal{C}$ a perturbation region of interest. Fix confidence and coverage parameter $\psi,R\in(0,1)$. Fix a layer $i\in\{1,\dots,N-2\},$ and for each $j=1,\dots,d_{i}$ compute an estimated reachable set $[\overline{l}_{j},\underline{u}_{j}]$ for neuron $z^{(i)}_{j},$ using $n^{\prime}\geq m\frac{ln(1-\psi^{1/m})}{ln(1-(1-R)/m)}$ independently and uniformly sampled point from $\mathcal{C},$ as by Proposition 3.2. Let $\overline{{\bm{l}}}=(\overline{l}_{1},\dots,\overline{l}_{d_{i}})$ and $\underline{{\bm{u}}}=(\underline{u}_{1},\dots,\underline{u}_{d_{i}}).$ Let ${{\bm{A}}},{\bm{b}}$ be the vectors of the linear bounds coefficients and biases (inductively) computed for the ReLU layer $i+1$, and such that, with probability $\geq\displaystyle{\psi^{\frac{d_{i+1}+d_{i+2}+\cdots+d_{N}}{m}},}$ for any $n\in\mathbb{N}$ in any sequence of $n$ points uniformly and independently sampled from $\mathcal{C},$ for at least $n\times\left(1-\frac{1-R}{m}\sum_{j=i+1}^{N}d_{j}\right)$ points ${\bm{x}}$ in $X$ it holds that
@@
where ${\bm{v}}_{{\bm{x}}}$ is the vector of pre-activation values of the neurons in the layer $i$ when the input is ${\bm{x}}.$
Then, with probability $\geq\displaystyle{\psi^{\frac{d_{i}}{m}},}$ it holds that for any $n\in\mathbb{N}$ in any sequence $X$ of $n$ points uniformly and independently sampled from $\mathcal{C},$ for at least $n\times\left(1-\frac{1-R}{m}d_{i}\right)$ points ${\bm{x}}$ in $X$ it holds that:

### v2

@@ 2.3 Linear Relaxation-based Perturbation Analysis (LiRPA) Approaches
For a DNN composed of $N$ layers, the base case is defined as $\underline{{\bm{A}}}^{(N)}=I$. Moreover, in the case of a single output node, $\underline{{\bm{A}}}^{(N-1)}={\bm{w}}^{\top}$, where $\bm{w}$} is the weight vector of the final layer. For the remaining layers $i\in\{N-2,\dots,1\}$, the linear relaxation is propagated using the recurrence $\underline{{\bm{A}}}^{(i)}=\underline{{\bm{A}}}^{(i+1)}\,\underline{{\bm{D}}}^{(i)}\,{\bm{W}}^{(i)}$, where ${\bm{W}}^{(i)}$ is the inter level weight matrix and $\underline{{\bm{D}}}^{(i)}$ is a diagonal matrix encoding the linear relaxation of the activation function at the $i$-th layer. For each layer $i$, we also recursively compute a bias vector $\underline{{\bm{b}}}^{(i)}$ based on the reachable sets of the $i$-th layer nodes and the matrix $\underline{{\bm{A}}}^{(i+1)}$. Each diagonal entry $\underline{{\bm{D}}}^{(i)}_{j,j}$ depends on the preactivation bounds of neuron $j$ (computed during the forward pass) and the sign of $\underline{{\bm{A}}}_{j}^{(i+1)}$, i.e., the coefficient associated with neuron $j$ in the linear relaxation of the next layer, for which we report the explicit formulas used for computing $\underline{{\bm{D}}}^{(i)}$ and ${\bm{b}}^{(i)}$ in the example provided below. At the end of the process, a provable lower bound on the minimum of $f$ in $\mathcal{C}$ (the $\ell_{\infty}$ norm ball around ${\bm{x}}_{0}$) is then easily obtained using Hölder’s inequality (crown) as
[…]
with ${\bm{c}}_{\text{LiRPA}}=\underline{{\bm{A}}}^{(N)}\underline{{\bm{b}}}^{(N-1)}+\ldots+\underline{{\bm{A}}}^{(2)}\underline{{\bm{b}}}^{(1)}$.
To provide the reader a concrete and practical illustration of this approach, in the following, we present a simple example of linear bound computation for the toy DNN shown in Figure 2, using CROWN, a state-of-the-art LiRPA-based method (crown).
@@
$x_{1}$$x_{2}$$z_{1}^{(1)}$$z_{2}^{(1)}$$\hat{z}_{1}^{(1)}$$\hat{z}_{2}^{(1)}$$z_{1}^{(2)}$$z_{2}^{(2)}$$\hat{z}_{1}^{(2)}$$\hat{z}_{2}^{(2)}$f(x)f(x)223-31144ReLUReLU44222-211ReLUReLU2-211 [2,2][-2,2] [1,3][-1,3] ${\bm{z}}^{(1)}$ [5,7][-5,7] [10,18][-10,18] $[0,18]$ $\hat{{\bm{z}}}^{(1)}$ $[0,7]$ ${\bm{z}}^{(2)}$ [36,28][-36,28] $[0,32]$ $\hat{{\bm{z}}}^{(2)}$ $[0,28]$ $[0,32]$ $\bm{[-33,18.9]}$ [56,32][-56,32] [42,24.3][-42,24.3] $z^{(3)}$
Figure 2: Toy DNN used in this example. Intervals reported in green are the exact output reachable set computed via MIP, in black are the ones of IBP, and finally, in purple, the results for CROWN considering the input [[2,2],[1,3]][[-2,2],[-1,3]].
Consider a neural network with two inputs, two hidden layers with ReLU activation, and one single output. Following the notation introduced in Sec. 2 we define:
@@
By propagating these intervals through the DNN, we obtain the interval [56,32][-56,32] as the output reachable set. Given the reasonable size of the neural network, before computing the linear lower and upper bounds using LiRPA, we employed the sound and complete MIP (MIP) solver to compute the true min and max of the function, respectively, which correspond to [33,18.89][-33,18.89], highlighted in green in Fig. 2.
To compute the lower and upper bounds using CROWN (crown), we employ LiRPA’s backward computation strategy. To this end, it is helpful to represent the neural network as reported in Fig. 3.
@@
We note that $\hat{{\bm{z}}}^{(2)}$ and $\hat{{\bm{z}}}^{(1)}$ contain non-linear activation functions (ReLU), and we have to linearize them to keep the linear relationship between the output and these hidden layers. To this end, we create $2\times\#$ReLU layers (for the lower and upper bound, respectively) diagonal matrices $\underline{{\bm{D}}}^{(2)},\overline{{\bm{D}}}^{(2)},\underline{{\bm{D}}}^{(1)},\overline{{\bm{D}}}^{(1)}$ and bias vectors $\underline{{\bm{b}}}^{(2)},\overline{{\bm{b}}}^{(2)},\underline{{\bm{b}}}^{(1)},\overline{{\bm{b}}}^{(1)}$ reflecting the impact of each ReLU nodes on the final output. We report for simplicity here the original definition provided in (crown) (a similar definition is applied to compute the $i$-th layer $\overline{{\bm{D}}}^{(i)}$ and $\overline{{\bm{b}}}^{(i)}$):
[…]
@@
$\geq\underline{{\bm{A}}}^{(3)}\underline{{\bm{D}}}^{(2)}z^{(2)}({\bm{x}})$ $\text{computing a linearization for }\hat{z}^{(2)}$
$\geq\underbrace{\underline{{\bm{A}}}^{(3)}\underline{{\bm{D}}}^{(2)}{\bm{W}}^{(2)}}_{\underline{{\bm{A}}}^{(2)}}\hat{z}^{(1)}({\bm{x}})$ $\text{rewriting }{z}^{(2)}={\bm{W}}^{(2)}\hat{z}^{(1)}$
$\geq\underline{{\bm{A}}}^{(2)}\underline{{\bm{D}}}^{(1)}z^{(1)}({\bm{x}})$ $\text{computing a linear bound for }\hat{z}^{(1)}$
$\geq\underbrace{\underline{{\bm{A}}}^{(2)}\underline{{\bm{D}}}^{(1)}{\bm{W}}^{(1)}}_{\underline{{\bm{A}}}^{(1)}}({\bm{x}})$ $\text{rewriting }{z}^{(1)}={\bm{W}}^{(1)}\hat{z}^{(0)}={\bm{W}}^{(1)}({\bm{x}})$
@@
Finally, we compute the sum in the bias vectors $\underline{d}=\underline{{\bm{A}}}^{(3)}\underline{{\bm{b}}}^{(2)}+\underline{{\bm{A}}}^{(2)}\underline{{\bm{b}}}^{(1)}=-35.88$ and $\overline{d}=\overline{{\bm{A}}}^{(3)}\overline{{\bm{b}}}^{(2)}+\overline{{\bm{A}}}^{(2)}\overline{{\bm{b}}}^{(1)}=12.27$. The final linear relation is thus $\underline{f}({\bm{x}})\geq\underline{{\bm{A}}}^{(1)}({\bm{x}})+\underline{d}$ and $\overline{f}({\bm{x}})\leq\overline{{\bm{A}}}^{(1)}({\bm{x}})+\overline{d}$. Using Hölder’s inequality (crown), we obtain
[…]
@@
[…]
@@
As we can notice, we obtain a tight over-approximation of the true lower and upper bounds, significantly improving the bounds derived with naive IBP ([-56,32]).
@@ 3 Probabilistically Tightened LiRPA via Underestimation
In this section, we theoretically investigate whether and how it is possible to compute tight intermediate reachable sets, which directly impact the linear output bound computation. As highlighted in Alg. 3, the entire linearization process crucially depends on the bounds computed at line 3. However, computing exact values for such bounds is generally infeasible, as the problem has been shown to be NP-hard Reluplex; fastlin. Motivated by the speculation of acrown, we therefore explore efficient alternatives for approximating the exact (unknown) values of these intermediate bounds as tightly as possible. To this end, we study the impact of a sampling-based approach and what type of probabilistic guarantees can be achieved with it. In detail, we begin employing the statistical prediction of tolerance limits results wilks, which allows a closed-form derivation of the required sample size to achieve a desired confidence level. However, as we will show, this method only quantifies the fraction of the perturbation region where the guarantees may fail, without addressing the magnitude of potential violations relative to the estimated bounds. Consequently, the resulting probabilistic certificates are weaker than those provided by related approaches such as (proven; randomizedSmoothing), where guarantees hold across the entire perturbation region. To mitigate this limitation, we first introduce a qualitative extension of Wilks’ guarantees adapted to our setting. While novel, this bound can become overly loose in high-dimensional scenarios. To address this issue, we then develop a new theoretical and practical result based on extreme value theory haan2006extreme, which allows us to tightly bound the magnitude of possible violations with respect to the estimated bounds.
We now present all the theoretical and practical components to compute tightened bounds with probabilistic guarantees within our PT-LiRPA approach. Our approach is based on a statistical methodology that allows us to compute estimates on the neural network’s output in the form $\min_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})$ and $\max_{{\bm{x}}\in\mathcal{C}}f({\bm{x}})$, which hold with high confidence. In particular, we employ the tools of statistical prediction of tolerance limits of (wilks). Fix a node $z^{(i)}_{j}$ in the neural network and, as before, let $z^{(i)}_{j}({\bm{x}})$ be the preactivation value for the node when the input to the network is a vector $\bm{x}$}. Let ${\bm{x}}_{1},\dots\bm{x}_{n}$ be $n$ points (vectors) independently and uniformly sampled from the perturbation set of interest $\mathcal{C}.$ We compute $z^{(i)}_{j}$’s (estimated) pre-activated bounds as:
@@
However, based on the results of (wilks), for any $\psi,R\in(0,1)$, we can choose the sample size $n$ that guarantees that the estimated reachable set is correct with probability $\psi$ for at least a fraction $R$ of points in the perturbation set $\mathcal{C}$. Crucially, this statistical result does not require any knowledge of the probability distribution governing the output of our function of interest. Formally, we have the following.
@@
Let $n$ the number of samples employed in the computation and the interval $\overline{l}^{(i)}_{j},\underline{u}^{(i)}_{j}$], where $\overline{l}_{j}^{(i)}$ and $\underline{u}_{j}^{(i)}$ are the minimum and maximum pre-activation values observed in the sample, respectively. Fix $R\in(0,1)$, then for any further possibly infinite sequence of samples from $\mathcal{C}$, the probability that $\overline{l}_{j}^{(i)},\underline{u}_{j}^{(i)}]$ is correct [^3] In the sense that there exists $C^{\prime}\subseteq C$ such that $|C^{\prime}|/|C|\geq R$ and for all ${\bm{x}}\in C^{\prime}$ it holds that $z^{(i)}_{j}({\bm{x}})\in[\overline{l}_{j}^{(i)},\underline{u}_{j}^{(i)}].$ for at least a fraction $R$ of points is at least $\psi=n\cdot\int_{R}^{1}x^{n-1}\;dx=(1-R^{n})$.
Hence, following Lemma 3.1, whose proof follows from (wilks), for any desired confidence level $\psi$, and lower bound fraction $R$, we can compute the number $n$ of samples sufficient to obtain the provable probabilistic guarantees on the desired reachable set accuracy. Specifically, we have that if we use a sample of $n\geq\frac{\ln{(1-\psi)}}{\ln{(R)}}$ input points and with them we obtain an estimated reachable set $\overline{l}^{(i)}_{j},\underline{u}^{(i)}_{j}$], then with probability $\psi$ at most a fraction (1R)(1-R) of points in an indefinitely larger future sample could fall outside such reachable set. By taking into account the total number of neurons in the DNN, and using a independently chosen set of points for each neuron, we can extend the above result to compute reachable sets for each neuron that are simultaneously correct with probability $\psi$ for at least a fraction $R$ of the points in $\mathcal{C}$:
@@
We will now show that we can combine the (probabilistically valid) estimation computed on the reachable sets of individual neurons using a LiRPA approach, so as to obtain, first, a probabilistically valid over-approximation of any ReLU layer in the DNN and thus on the final network’s lower and upper outputs. In detail, we begin by proving that the estimated reachable sets used to produce the vectors $\bm{A}$} and $\bm{b}$}, together with the linearization applied to each ReLU layer of the network (as in Sec. 2.3), yield a probabilistically sound over-approximation that covers at least a fraction $R$ of the perturbation region $\mathcal{C}$.
@@
Consider a ReLU DNN with $m$ neurons distributed over $N$ layers, and $\mathcal{C}$ a perturbation region of interest. Fix confidence and coverage parameter $\psi,R\in(0,1)$. Fix a layer $i\in\{1,\dots,N-2\},$ and for each $j=1,\dots,d_{i}$ compute an estimated reachable set $\overline{l}_{j},\underline{u}_{j}]$ for neuron $z^{(i)}_{j}$, using $n^{\prime}\geq m\frac{ln(1-\psi^{1/m})}{ln(1-(1-R)/m)}$ independently and uniformly sampled point from $\mathcal{C},$ as by Proposition 3.2. Let $\overline{{\bm{l}}}=(\overline{l}_{1},\dots,\overline{l}_{d_{i}})$ and $\underline{{\bm{u}}}=(\underline{u}_{1},\dots,\underline{u}_{d_{i}}).$ Let ${{\bm{A}}},{\bm{b}}$ be the vectors of the linear bounds coefficients and biases (inductively) computed for the ReLU layer $i+1$, and such that, with probability $\geq\displaystyle{\psi^{\frac{d_{i+1}+d_{i+2}+\cdots+d_{N}}{m}},}$ for any $n\in\mathbb{N}$ in any sequence of $n$ points uniformly and independently sampled from $\mathcal{C},$ for at least $n\times\left(1-\frac{1-R}{m}\sum_{j=i+1}^{N}d_{j}\right)$ points $\bm{x}$} in $X$ it holds that
@@
where ${\bm{v}}_{{\bm{x}}}$ is the vector of pre-activation values of the neurons in the layer $i$ when the input is $\bm{x}$}. Then,
  1. 1.
    with probability $\geq\displaystyle{\psi^{\frac{d_{i}}{m}},}$ it holds that for any $n\in\mathbb{N}$ in any sequence $X$ of $n$ points uniformly and independently sampled from $\mathcal{C},$ for at least $n\times\left(1-\frac{1-R}{m}d_{i}\right)$ points $\bm{x}$} in $X$ it holds that: